# Rectifying Regression in Reinforcement Learning

### v1

@@ 1 Introduction
Value-based methods are ubiquitous in reinforcement learning (RL) (Sutton & Barto, 2018) and contextual bandits (Lattimore & Szepesvári, 2020), where the goal is to predict rewards and then choose actions that maximize expected returns. The “natural objective function” (Chapter 9.2 of (Sutton & Barto, 2018)) in these settings is the mean squared error (MSE). Traditionally in RL, value-based algorithms that leverage function approximation aim to minimize the MSE between a learned value function and observed returns, as popularized by algorithms such as $Q$-learning and its extension to function approximation (Ernst et al., 2005; Riedmiller, 2005; Mnih et al., 2015). Despite its success in many practical scenarios, recent theoretical (Foster & Krishnamurthy, 2021; Ayoub et al., 2024) and empirical (Farebrother et al., 2024) findings highlight that minimizing MSE on the data can yield worse decision-making performance (in terms of expected returns) than alternative empirical losses. In particular, minimizing alternative losses can achieve faster learning when the optimal value $v^{\star}$ is close to the maximum possible value.
A promising alternative is the log-loss (cross-entropy loss), which we argue aligns more closely with the mean absolute error (MAE)—a tighter surrogate for decision quality. Indeed, it has been shown that, under certain problem structures, controlling the MAE can lead to faster convergence (Foster & Krishnamurthy, 2021) and a smaller suboptimality gap (Ayoub et al., 2024) than controlling MSE. Building on these insights, we analyze value-based methods that employ log-loss and demonstrate their advantages over those using squared loss. Our exposition initially restricts attention to the offline contextual bandit (or reward-sensitive classification (Elkan, 2001)) setting, where one aims to learn a policy from a fixed dataset of context–reward pairs. However, the core ideas and proofs naturally extend to more general RL problems with function approximation (Antos et al., 2007; Chen & Jiang, 2019; Ayoub et al., 2024).
We propose a reparameterized version of the categorical cross-entropy loss ($\ell_{\operatorname{cat}}$), which learns a multi-category distribution over possible outcomes while still accurately recovering the mean of a bounded random variable as was similarly done by Lyle et al. (2019). This addresses the irreducible bias found in purely bin-based approaches.
@@ 2.3 Solution: Value-Based Methods for Classification
However, directly optimizing this empirical risk is generally computationally intractable (NP hard), even for relatively simple policy classes (Ben-David et al., 2003; Feldman et al., 2012). This motivates the use of value-based methods for classification, which leverage a class of candidate value functions to reformulate the problem as a regression task—often solvable via gradient descent.
@@ 2.4 The Choice of the Loss Function
In the context of reinforcement learning, value-based methods commonly minimize the squared loss (Sutton & Barto, 2018; Szepesvári, 2022):
@@ 3.1 Positive Results
We highlight that minimizing the log-loss (i.e., $\ell_{\operatorname{log}}$) yields bounds that scale with $(1-v^{\star})$, which can be small in problems where the optimal policy achieves a reliable goal or accumulates near-maximal return. The following lemma adapts the result of Foster & Krishnamurthy (2021) to the rewards-based setting.
@@
The proof of Lemma˜3.1 can be found in Appendix˜B. This result first appeared for cost-sensitive classification (Foster & Krishnamurthy, 2021) and was extended to batch RL (with costs) in Ayoub et al. (2024). By adapting their proofs, one obtains the same bound for batch RL with rewards. Observe that their analysis also proceeds by reducing the classification (or RL) objective to bounding the MAE between $\hat{f}_{\log}$ and $r$. We now show that, under certain conditions ($v^{\star}\approx 1$), one can achieve small MAE while the rMSE remains large.
@@ 3.2 Negative Results
In this section, we argue that rMSE is not the most “natural objective function” (Chapter 9.2 of (Sutton & Barto, 2018)) for RL. We make two claims. The first claim (Proposition˜3.2) is that there are problems where the rMSE of $\hat{f}_{\log}$ is at least $\tfrac{1}{\sqrt{n}}$, while its MAE is as small as $\tfrac{1}{n}$. The second claim (Lemma˜3.3) is that if one directly minimizes the squared loss $\ell_{\operatorname{sq}}$, then there exist problems for which the resulting estimator suffers a large MAE (and hence rMSE) of order $1/\sqrt{n}$, even when $v^{\star}\approx 1$.
@@
Finally, we show that this “over-weighting of rare events” carries over to minimizing the squared loss directly, as commonly done by value-based RL algorithms with function approximation (Sutton & Barto, 2018).
@@ 4 Reparameterizing the Categorical Cross Entropy Loss
Recall that the canonical categorical cross-entropy (cat-loss) can be written as the negative log-likelihood of an exponential family (Brown, 1987). Let $y\in[0,1]$ be a scalar and $\theta\in\mathbb{R}^{K}$. In its canonical form, the cat-loss can be (naively) used for value learning in reinforcement learning,
@@
Since the cat-loss is a generalization of the log-loss, a first-order bound, similar to that of Lemma˜3.1, can be shown for the cat-loss. Thus the cat-loss also does well when the optimal return $v^{\star}\approx 1$, albeit with an additional factor of $K$ scaling the bound due to concentration arguments (Zhang, 2006; Grünwald & Mehta, 2020).
@@
This mean-preserving strategy can be generalized to other distribution-based losses derived from exponential families. For instance, HL Gauss (Imani & White, 2018), which has been shown to have very low test MAE (Imani et al., 2024), can be reparameterized similarly to retain its distributional modeling benefits for classification, while accurately recovering the mean of continuous targets. In a broader sense, any exponential-family log-likelihood can be adapted for dual “classification”-regression usage by carefully selecting sufficient statistics that embed the raw target $y$, thus allowing us to harness the benefits of classification [^1] Cross-entropy losses losses (i.e., categorical cross entropy) for reinforcement learning and continue regressing (Farebrother et al., 2024).
@@ 5 Numerical Experiments
We evaluate fitted Q-iteration (Ernst et al., 2005; Szepesvári, 2022) trained with squared loss (Square), log-loss (Log), and cat-loss (Cat) on the inverted pendulum environment (Lagoudakis & Parr, 2003; Riedmiller, 2005), where the goal is to keep an inverted pendulum balanced by applying the correct forces. The state space is two-dimensional (angle and angular velocity), and there are three actions (left, right, do nothing). The environment dynamics follow Lagoudakis & Parr (2003), with two modifications: (i) when the pendulum falls below horizontal, the state terminates, and (ii) the angular momentum is clipped to [-5,5] to facilitate the use of Fourier features (Konidaris et al., 2011) of orders 2 and 3. The agent receives a reward of 0 for staying upright and -1 for falling, with a discount factor of $\gamma=0.99$. All datasets are collected by a policy that selects actions uniformly at random until failure (which typically occurs after 6 steps). We then evaluate whether the learned policies keep the inverted pendulum above horizontal after 3000 steps and report the policy’s failure rate.
@@
For minimizing Square, we use its closed-form solution; for minimizing Log, we apply Newton’s method (Sun & Tran-Dinh, 2019); and for minimizing Cat, we use a limited-memory BFGS method (Schmidt et al., 2009). All three methods are guaranteed to converge to their respective optima superlinearly.
@@ Appendix A Additional Notation
The second part of the lemma follows from the argument in the proof of Lemma 1 in Foster & Krishnamurthy (2021), where bounding the mean absolute error appears as an intermediate step. ∎
@@ Appendix D Technical Results
Lemma D.1 (Lemma 1 of Foster & Krishnamurthy (2021)).
@@
Lemma D.2 (Lemma A.1 of Ayoub et al. (2024)).
@@ D.1 Concentration for the Log-Loss Estimator
Foster & Krishnamurthy (2021) establish the following concentration result for $\hat{f}_{\log}$. We restate it here for completeness.
@@
The result follows directly from the last equation on page 24 of the arXiv version of Foster & Krishnamurthy (2021), taking $A=1$. ∎

### v2

@@ 1 Introduction
Value-based methods are ubiquitous in reinforcement learning (RL) (sutton2018reinforcement) and contextual bandits (lattimore2020bandit), where the goal is to predict rewards and then choose actions that maximize expected returns. The “natural objective function” (Chapter 9.2 of (sutton2018reinforcement)) in these settings is the mean squared error (MSE). Traditionally in RL, value-based algorithms that leverage function approximation aim to minimize the MSE between a learned value function and observed returns, as popularized by algorithms such as $Q$-learning and its extension to function approximation (ernst2005tree; riedmiller2005neural; Mnih2015HumanlevelCT). Despite its success in many practical scenarios, recent theoretical (foster2021efficient; ayoub2024switching) and empirical (farebrother2024stop) findings highlight that minimizing MSE on the data can yield worse decision-making performance (in terms of expected returns) than alternative empirical losses. In particular, minimizing alternative losses can achieve faster learning when the optimal value $v^{\star}$ is close to the maximum possible value.
A promising alternative is the log-loss (cross-entropy loss), which we argue aligns more closely with the mean absolute error (MAE)—a tighter surrogate for decision quality. Indeed, it has been shown that, under certain problem structures, controlling the MAE can lead to faster convergence (foster2021efficient) and a smaller suboptimality gap (ayoub2024switching) than controlling MSE. Building on these insights, we analyze value-based methods that employ log-loss and demonstrate their advantages over those using squared loss. Our exposition initially restricts attention to the offline contextual bandit (or reward-sensitive classification (elkan2001foundations)) setting, where one aims to learn a policy from a fixed dataset of context–reward pairs. However, the core ideas and proofs naturally extend to more general RL problems with function approximation (antos2007fitted; chen2019information; ayoub2024switching).
We propose a reparameterized version of the categorical cross-entropy loss ($\ell_{\operatorname{cat}}$), which learns a multi-category distribution over possible outcomes while still accurately recovering the mean of a bounded random variable as was similarly done by lyle2019comparative. This addresses the irreducible bias found in purely bin-based approaches.
@@ 2.3 Solution: Value-Based Methods for Classification
However, directly optimizing this empirical risk is generally computationally intractable (NP hard), even for relatively simple policy classes (ben2003difficulty; feldman2012agnostic). This motivates the use of value-based methods for classification, which leverage a class of candidate value functions to reformulate the problem as a regression task—often solvable via gradient descent.
@@ 2.4 The Choice of the Loss Function
In the context of reinforcement learning, value-based methods commonly minimize the squared loss (sutton2018reinforcement; szepesvari2022algorithms):
@@ 3.1 Positive Results
We highlight that minimizing the log-loss (i.e., $\ell_{\operatorname{log}}$) yields bounds that scale with $(1-v^{\star})$, which can be small in problems where the optimal policy achieves a reliable goal or accumulates near-maximal return. The following lemma adapts the result of foster2021efficient to the rewards-based setting.
@@
The proof of Lemma˜3.1 can be found in Appendix˜B. This result first appeared for cost-sensitive classification (foster2021efficient) and was extended to batch RL (with costs) in ayoub2024switching. By adapting their proofs, one obtains the same bound for batch RL with rewards. Observe that their analysis also proceeds by reducing the classification (or RL) objective to bounding the MAE between $\hat{f}_{\log}$ and $r$. We now show that, under certain conditions ($v^{\star}\approx 1$), one can achieve small MAE while the rMSE remains large.
@@ 3.2 Negative Results
In this section, we argue that rMSE is not the most “natural objective function” (Chapter 9.2 of (sutton2018reinforcement)) for RL. We make two claims. The first claim (Proposition˜3.2) is that there are problems where the rMSE of $\hat{f}_{\log}$ is at least $\tfrac{1}{\sqrt{n}}$, while its MAE is as small as $\tfrac{1}{n}$. The second claim (Lemma˜3.3) is that if one directly minimizes the squared loss $\ell_{\operatorname{sq}}$, then there exist problems for which the resulting estimator suffers a large MAE (and hence rMSE) of order $1/\sqrt{n}$, even when $v^{\star}\approx 1$.
@@
Finally, we show that this “over-weighting of rare events” carries over to minimizing the squared loss directly, as commonly done by value-based RL algorithms with function approximation (sutton2018reinforcement).
@@ 4 Reparameterizing the Categorical Cross Entropy Loss
Recall that the canonical categorical cross-entropy (cat-loss) can be written as the negative log-likelihood of an exponential family (brown1987fundamentals). Let $y\in[0,1]$ be a scalar and $\theta\in\mathbb{R}^{K}$. In its canonical form, the cat-loss can be (naively) used for value learning in reinforcement learning,
@@
Since the cat-loss is a generalization of the log-loss, a first-order bound, similar to that of Lemma˜3.1, can be shown for the cat-loss. Thus the cat-loss also does well when the optimal return $v^{\star}\approx 1$, albeit with an additional factor of $K$ scaling the bound due to concentration arguments (zhang2006varepsilon; grunwald2020fast).
@@
This mean-preserving strategy can be generalized to other distribution-based losses derived from exponential families. For instance, HL Gauss (imani2018improving), which has been shown to have very low test MAE (imani2024investigating), can be reparameterized similarly to retain its distributional modeling benefits for classification, while accurately recovering the mean of continuous targets. In a broader sense, any exponential-family log-likelihood can be adapted for dual “classification”-regression usage by carefully selecting sufficient statistics that embed the raw target $y$, thus allowing us to harness the benefits of classification [^1] Cross-entropy losses losses (i.e., categorical cross entropy) for reinforcement learning and continue regressing (farebrother2024stop).
@@ 5 Numerical Experiments
We evaluate fitted Q-iteration (ernst2005tree; szepesvari2022algorithms) trained with squared loss (Square), log-loss (Log), and cat-loss (Cat) on the inverted pendulum environment (lagoudakis2003least; riedmiller2005neural), where the goal is to keep an inverted pendulum balanced by applying the correct forces. The state space is two-dimensional (angle and angular velocity), and there are three actions (left, right, do nothing). The environment dynamics follow lagoudakis2003least, with two modifications: (i) when the pendulum falls below horizontal, the state terminates, and (ii) the angular momentum is clipped to [-5,5] to facilitate the use of Fourier features (konidaris2011value) of orders 2 and 3. The agent receives a reward of 0 for staying upright and -1 for falling, with a discount factor of $\gamma=0.99$. All datasets are collected by a policy that selects actions uniformly at random until failure (which typically occurs after 6 steps). We then evaluate whether the learned policies keep the inverted pendulum above horizontal after 3000 steps and report the policy’s failure rate.
@@
For minimizing Square, we use its closed-form solution; for minimizing Log, we apply Newton’s method (sun2019generalized); and for minimizing Cat, we use a limited-memory BFGS method (pmlr-v5-schmidt09a). All three methods are guaranteed to converge to their respective optima superlinearly.
@@ Appendix A Additional Notation
The second part of the lemma follows from the argument in the proof of Lemma 1 in foster2021efficient, where bounding the mean absolute error appears as an intermediate step. ∎
@@ Appendix D Technical Results
Lemma D.1 (Lemma 1 of foster2021efficient).
@@
Lemma D.2 (Lemma A.1 of ayoub2024switching).
@@ D.1 Concentration for the Log-Loss Estimator
foster2021efficient establish the following concentration result for $\hat{f}_{\log}$. We restate it here for completeness.
@@
The result follows directly from the last equation on page 24 of the arXiv version of foster2021efficient, taking $A=1$. ∎